A transformer has 200 primary windings ans 20 secondary windings. If the primary voltage is 10 V, find the secondary voltage.
Answer by rfadrogane(214)

A transformer has 200 primary windings and 20 secondary windings. If the primary voltage is 10 V, find the secondary voltage.
Sol'n:
Let: N1 - # of turns primary winding
N2 - # of turns secondary winding
V1 - voltage on primary winding
V2 - voltage on secondary winding
N1/N2 = V1/V2
V2 = V1(N2/N1)
V2 = 10(20/200)
V2 = 1 V ----ans.